# FlowMalTrans: Unsupervised Binary Code Translation for Malware Detection Using Flow-Adapter Architecture

### v1

@@ 5.3 Testing on Binary Code Translation
We use SacreBLEU sacrebleu to compute the BLEU score as the average precision of unigrams, bigrams, trigrams, and 4-grams.
Testing Datasets. We use zlib, coreutils, and diffutils for testing. Note that these programs are not used for training FlowMalTrans.
@@
We compare FlowMalTrans to the baseline UnsuperBinTrans. Since UnsuperBinTrans  focuses solely on two ISAs (X86-64 and ARM32), we train UnsuperBinTrans on the same training datasets for additional ISA pairs. We can see that FlowMalTrans outperforms UnsuperBinTrans across all ISA pairs.
Table 3: Malware detection results.
@@
Train & Test on Optimal
@@ 5.4 Malware Detection Task
Malware Detection Model. We use the LSTM model lstm-2018 to detect malware, designed as two layers. The model details are provided in Appendix A.5. We first extract the token embeddings from FlowMalTrans and integrate them into the input layer of LSTM. As a result, when a binary is fed into LSTM, each input token is represented as its embedding. To further enrich the task, we also apply a CNN model, with the results presented in Appendix A.6. In the following, we focus on the results related to LSTM.
Task-Specific Training Datasets. We collect $2140$, $1740$, $1581$, $1430$, $962$, $545$, and $128$ malware samples from VirusShare.com VirusShare for X86-64, i386, ARM32, MIPS32, PPC32, M68K, and ARM64, respectively. We spend significant efforts in collecting malware, especially for PPC32, M68K, and ARM64. Notably, our approach only needs malware in a high-resource ISA to train a malware detection model, significantly reducing the data collection burden for low-resource ISAs.
@@ 5.5 Model Comparison
Comparison with Baseline Methods. The first baseline is UnsuperBinTrans. We use it to translate binaries from the other ISAs to X86-64 and use LSTM to test the translated binaries. The results in Table 3(a) show that FlowMalTrans achieves better translation quality, leading to improved malware detection performance.
The second baseline is UniMap wang-security23, which covers four ISAs (X86-64, ARM32, MIPS32, and PPC32). To ensure a comprehensive comparison, we train it on the same datasets to cover the three additional ISAs (i386, ARM64 and M68K). Table 3(a) show that FlowMalTrans outperforms UniMap across all ISA pairs.
The third baseline is CrossIns2Vec wang2024learning, which supports four ISAs (X86-64, ARM32, MIPS32, and PPC32). As it is not open-sourced, the comparison is limited to these four ISAs. The results show that FlowMalTrans achieves better performance than CrossIns2Vec.
In the fourth baseline analyzes IR code. We assess whether a model trained on X86-64 IR code can be reused to test IR code in other ISAs. We use angr vex-ir to lift binaries into IR. We train LSTM using the same task-specific training dataset in X86-64, and apply the model to test the same task-specific datasets in other ISAs. The results indicate IR alone does not magically allow a model trained on one ISA to be reused for other ISAs. (more explanation is discussed in Appendix A.10).
Comparison with Optimal Model. The results of the Same-ISA model are shown in Table 3(b). Comparing the results with our model in Table 3(a), we observe: (1) our model achieves performance close to Same-ISA model for i386, ARM32 and MIPS32; (2) our model significantly outperforms Same-ISA model for ARM64, PPC32 and M68K.
For i386, ARM32, and MIPS32, the results align with expectations: the Same-ISA model, trained and tested on the same ISA, outperforms FlowMalTrans, which is trained on X86-64 and tested on other ISAs (via translation). However, for ARM64, PPC32, and M68K, our model outperforms the Same-ISA model, due to the insufficient malware samples used to train the Same-ISA model, highlighting the importance of sufficient training data to achieve desirable performance. While including more data could enhance the model’s performance, collecting malware samples for low-resource ISAs can be challenging. Our approach—translating binaries to x86-64—addresses data scarcity effectively.
@@ 6 Conclusion
Malware Packing and Obfuscation. The practice of malware packing entails hiding malicious code inside files that appear benign. Our research excludes malware packing from consideration, as it lies beyond our analytical scope. Our investigation centers on examining how well malware detection models transfer across different instruction set architectures (ISAs) through binary translation. The malware specimens employed in our detection experiments are uncompressed, enabling their examination and reverse engineering using analytical tools such as IDA Pro IDA.
Additionally, when dealing with compressed malware, one approach involves utilizing sophisticated decompression utilities, including PEiD peid and OllyDbg OllyDbg, to initially decompress the malware before analyzing the resulting content.
Code obfuscation methods present significant obstacles for malware identification systems. The malware specimens utilized in our research originated from VirusShare.com VirusShare, a collection platform that gathers malware encountered in real-world environments. It is commonly understood that such malware frequently implements obfuscation strategies to circumvent detection systems. Nevertheless, we do not possess definitive information about the particular obfuscation methods employed in each malware sample, which complicates evaluating resistance against specific obfuscation approaches.
@@
For evaluation within the security domain, we utilized malicious software specimens obtained from the VirusShare.com platform VirusShare, a well-established digital repository that serves the cybersecurity research community. We acknowledge the substantial ethical responsibilities accompanying the use of such security-sensitive materials. Our research protocols emphasized legitimate scientific inquiry, prevention of inadvertent threat proliferation, and careful attention to confidentiality and regulatory compliance.
@@ A.1 Model Training Details
Model Pretraining. Pretraining is a key ingredient of unsupervised neural machine translation. Studies have shown that the pre-trained cross-lingual word embeddings which are used to initialize the lookup table, have a significant impact on the performance of an unsupervised machine translation model conneau2020unsupervised; conneau2019cross. We adopt this and pre-train both the encoders and decoders of FlowMalTrans to bootstrap the iterative process of our binary translation model. We employ causal language modeling (CLM) and masked language modeling (MLM) objectives to train the encoder and decoder.
@@
Back Translation (BT). We adopt the back-translation approach feldman2020neural to train our model in a translation setting, as shown in Figure 2. As an example, given an input basic block in one ISA $N$, we use the model in the inference mode to translate it to the other ISA $M$ (i.e. applying the encoder of $N$ and the decoder of $M$). This way, we obtain a pseudo-parallel basic block pair. We then train the model to predict the original basic block (i.e. applying the encoder of $M$ and the decoder of $N$) from this synthetic translation. As training progresses and the model improves, it will produce better synthetic basic block pairs through backtranslation, which will serve to further improve the model in the subsequent iterations. In BT, the latent code transformation is performed twice and trained along with the encoders/decoders; taking BT for the source ISA as an example: first in the source-to-target direction, then in the target-to-source direction as shown in the steps of 3 and 4 in Figure 2.
@@ A.9 Hyperparameter and Ablation Study
Normalizing Flows. We conduct an experiment to analyze the impact of normalizing flows on translation capability based on the following aspects: (1) exploring different flow-adapter architectures, such as Glow diederik2014adam; (2) varying the number of sequential flows to build the source/target flow; and (3) removing the flow-adapter architecture in FlowMalTrans.
@@ A.12 Uniqueness of Our Work
We propose an entirely different approach compared to the prior works wang2023can; wang2024learning, such as UniMap wang2023can, which also aims to reuse models across ISAs in order to resolve the data scarcity issue in low-resource ISAs.
@@
While UniMap learns cross-architecture instruction embeddings, our method focuses on translating binary code across ISAs. By translating code to a high-resource ISA, our approach offers several advantages. First, it allows the direct application of existing downstream models—trained on the high-resource ISA—to other ISAs through testing the translated code. In contrast, the works in wang2023can; wang2024learning require retraining the model using cross-architecture instruction embeddings. Furthermore, translating code from one ISA to another assists human analysts in understanding code from unfamiliar ISAs, supporting broader applications in code comprehension.
InnerEye zuo2018neural applies neural machine translation (NMT) techniques for binary code similarity comparison but does not perform binary code translation across ISAs. It uses two encoders from NMT models, where each encoder generates an embedding for a piece of binary code of a given pairs, and measures similarity based on embedding distance. In contrast, our approach focuses on translating binary code across ISAs.
Intermediate Representation (IR) is an architecture-agnostic representation that abstracts away various architectural differences among ISAs. However, despite this abstraction, significant variations still exist across IR code from different ISAs. Consequently, even when binaries from different ISAs—compiled from the same source code—are converted into a common IR, the resulting IR code often differs significantly, as discussed in ahmad-luo-2023-unsupervised. We compare malware detection performance across ISAs by translating binaries using FlowMalTrans against an approach that analyzes the IR code of binaries across ISAs. The results show our model enables superior cross-ISA malware detection (see the evaluation). An interesting direction is exploring translation based on IR code, which we leave for future work.

### v2

@@ 5.3 Testing on Binary Code Translation
We use the Bilingual Evaluation Understudy (BLEU) Papineni et al. (2002) to evaluate the translation performance, which is often used to evaluate the quality of machine-generated translations by measuring the $n$-gram overlap between the translation and reference. We set the tokenization of SacreBLEU Post (2018) to None, apply add-one smoothing, and use the default settings
Testing Datasets. We use three packages, zlib-1.2.11, coreutils-9.0, and diffutils-3.7, to test FlowMalTrans. Note that these packages are not included in the training dataset of FlowMalTrans. We compile them on the six ISAs using GCC-11.4.0 with different optimization levels (O0-O3) and use IDA Pro to disassemble them.
@@
We compare FlowMalTrans to the baseline UnsuperBinTrans. We use the open-source trained model of UnsuperBinTrans for this comparison. Note that UnsuperBinTrans  focuses solely on two ISAs (X86-64 and ARM32). To ensure a comprehensive comparison, we train UnsuperBinTrans on the same mono-architecture training datasets for the additional ISA pairs. We can see that FlowMalTrans outperforms UnsuperBinTrans across all ISA pairs and demonstrates satisfactory performance. Thus, FlowMalTrans has good translation quality and can effectively translate binaries across ISAs.
The average time to translate a basic block from one ISA to x86-64 is $10^{-4}$s, and the average number of blocks in a binary is $12,320$. Thus, translating a binary takes around $1.2$s. This demonstrates the good scalability of our approach.
Table 3: Malware detection results using LSTM. In Table (a), we compare the detection performance by translating malware using FlowMalTransUnsuperBinTrans. Then we evaluate it against UniMapCrossIns2Vecand the IR-based model. In Table (b), we give the results of the Same-ISA Model, which is trained and tested on the same ISA.
@@
Train & Test on Same-ISA
@@ 5.4 Malware Detection Task
Malware Detection Model. We use the LSTM model HaddadPajouh et al. (2018) to detect malware, designed as two layers. The model details are provided in Appendix A.5. We first extract the token embeddings from FlowMalTrans and integrate them into the input layer of LSTM. As a result, when a binary is fed into LSTM, each input token is represented as its embedding. To further enrich the task, we also apply a CNN model, with the results presented in Appendix A.6. In the following, we focus on the results related to LSTM.
Task-Specific Training Datasets. We collect $2140$, $1740$, $1581$, $1430$, $962$, $545$, and $128$ malware samples from VirusShare.com VirusShare (2020) for X86-64, i386, ARM32, MIPS32, PPC32, M68K, and ARM64, respectively. We spend significant efforts in collecting malware, especially for PPC32, M68K, and ARM64. Notably, our approach only needs malware in a high-resource ISA to train a malware detection model, significantly reducing the data collection burden for low-resource ISAs.
@@ 5.5 Model Comparison
Comparison with Baseline Methods. The first baseline is UnsuperBinTrans. As UnsuperBinTrans  focuses solely on two ISAs (X86-64 and ARM32), we train it on the same training datasets for the additional ISA pairs. We use it to translate binaries from the other ISAs to X86-64 and use LSTM to test the translated binaries. The results in Table 3(a) show that FlowMalTrans achieves better translation quality, leading to improved malware detection performance.
The second baseline is UniMap Wang et al. (2023a), which also aims to resolve the data scarcity issue in malware detection across ISAs. UniMap learns transferable knowledge for model reuse. It covers four ISAs, including X86-64, ARM32, MIPS32, and PPC32. To ensure a comprehensive comparison, we train it on the same datasets to cover the three additional ISAs (i386, ARM64 and M68K). Table 3(a) show that FlowMalTrans outperforms UniMap across all ISA pairs.
The third baseline is CrossIns2Vec Wang et al. (2024), which supports four ISAs (X86-64, ARM32, MIPS32, and PPC32). As it is not open-sourced, the comparison is limited to these four ISAs. The results show that FlowMalTrans achieves better performance than CrossIns2Vec.
The fourth baseline analyzes IR code. We assess whether a model trained on X86-64 IR code can be reused to test IR code in other ISAs. We use angr angr (2024) to lift binaries into IR. We train LSTM using the same task-specific training dataset in X86-64, and apply the model to test the same task-specific datasets in other ISAs. The AUC scores are $0.819$, $0.815$, $0.825$, $0.791$, $0.639$ and $0.476$ for i386, ARM32, ARM64, MIPS32, PPC32 and M68K, respectively. This indicates IR alone does not magically allow a model trained on one ISA to be reused for other ISAs. (more explanation is discussed in Appendix A.10).
Comparison with Same-ISA Model. The results of the Same-ISA model are shown in Table 3(b). When the LSTM model is trained and tested on the same ISA, it achieves AUC of $0.998$, $0.986$, $0.705$, $0.952$, $0.949$ and $0.790$ for i386, ARM32, ARM64, MIPS32, PPC32 and M68K, respectively. Comparing the results with our model in Table 3(a), we observe: (1) our model achieves performance close to Same-ISA model for i386, ARM32 and MIPS32; (2) our model significantly outperforms Same-ISA model for ARM64, PPC32 and M68K.
For i386, ARM32, and MIPS32, the results align with expectations: the Same-ISA model, trained and tested on the same ISA, outperforms FlowMalTrans, which is trained on X86-64 and tested on other ISAs (via translation). However, for ARM64, PPC32, and M68K, our model outperforms the Same-ISA model due to the insufficient malware samples used to train the Same-ISA model. For example, for M68K, only $80\%$ of $245$ malware samples are used for training. This highlights the importance of sufficient training data to achieve desirable performance. While including more data could enhance the model’s performance, collecting malware samples for low-resource ISAs can be challenging. Our approach—translating binaries to x86-64—addresses data scarcity effectively. In an extreme case, if only one binary in a given ISA is available, we can still detect whether it is malware using the X86-64-trained model.
@@ 6 Conclusion and Future Work
Large language Models (LLMs) show extraordinary capability in coding tasks and are popular utilized Xu et al. (2025); Ma et al. (2024). We leave using LLM for code translation as future work.
@@
Malware Packing and Obfuscation. The practice of malware packing entails hiding malicious code inside files that appear benign. Our research excludes malware packing from consideration, as it lies beyond our analytical scope. Our investigation centers on examining how well malware detection models transfer across different instruction set architectures (ISAs) through binary translation. The malware specimens employed in our detection experiments are uncompressed, enabling their examination and reverse engineering using analytical tools such as IDA Pro IDA (2023).
Additionally, when dealing with compressed malware, one approach involves utilizing sophisticated decompression utilities, including PEiD PEiD (2008) and OllyDbg OllyDbg (2000), to initially decompress the malware before analyzing the resulting content.
Code obfuscation methods present significant obstacles for malware identification systems. The malware specimens utilized in our research originated from VirusShare.com VirusShare (2020), a collection platform that gathers malware encountered in real-world environments. It is commonly understood that such malware frequently implements obfuscation strategies to circumvent detection systems. Nevertheless, we do not possess definitive information about the particular obfuscation methods employed in each malware sample, which complicates evaluating resistance against specific obfuscation approaches.
@@
For evaluation within the security domain, we utilized malicious software specimens obtained from the VirusShare.com platform VirusShare (2020), a well-established digital repository that serves the cybersecurity research community. We acknowledge the substantial ethical responsibilities accompanying the use of such security-sensitive materials. Our research protocols emphasized legitimate scientific inquiry, prevention of inadvertent threat proliferation, and careful attention to confidentiality and regulatory compliance.
@@ A.1 Model Training Details
Model Pretraining. Pretraining is a key ingredient of unsupervised neural machine translation. Studies have shown that the pre-trained cross-lingual word embeddings which are used to initialize the lookup table, have a significant impact on the performance of an unsupervised machine translation model Conneau et al. (2020); Conneau and Lample (2019). We adopt this and pre-train both the encoders and decoders of FlowMalTrans to bootstrap the iterative process of our binary translation model. We employ causal language modeling (CLM) and masked language modeling (MLM) objectives to train the encoder and decoder.
@@
Back Translation (BT). We adopt the back-translation approach Feldman and Coto-Solano (2020) to train our model in a translation setting, as shown in Figure 2. As an example, given an input basic block in one ISA $N$, we use the model in the inference mode to translate it to the other ISA $M$ (i.e. applying the encoder of $N$ and the decoder of $M$). This way, we obtain a pseudo-parallel basic block pair. We then train the model to predict the original basic block (i.e. applying the encoder of $M$ and the decoder of $N$) from this synthetic translation. As training progresses and the model improves, it will produce better synthetic basic block pairs through backtranslation, which will serve to further improve the model in the subsequent iterations. In BT, the latent code transformation is performed twice and trained along with the encoders/decoders; taking BT for the source ISA as an example: first in the source-to-target direction, then in the target-to-source direction as shown in the steps of 3 and 4 in Figure 2.
@@ A.9 Hyperparameter and Ablation Study
Normalizing Flows. We conduct an experiment to analyze the impact of normalizing flows on translation capability based on the following aspects: (1) exploring different flow-adapter architectures, such as Glow Diederik P. Kingma (2014); (2) varying the number of sequential flows to build the source/target flow; and (3) removing the flow-adapter architecture in FlowMalTrans.
@@ A.12 Uniqueness of Our Work
We propose an entirely different approach compared to the prior works Wang et al. (2023b, 2024), such as UniMap Wang et al. (2023b), which also aims to reuse models across ISAs in order to resolve the data scarcity issue in low-resource ISAs.
@@
While UniMap learns cross-architecture instruction embeddings, our method focuses on translating binary code across ISAs. By translating code to a high-resource ISA, our approach offers several advantages. First, it allows the direct application of existing downstream models—trained on the high-resource ISA—to other ISAs through testing the translated code. In contrast, the works in Wang et al. (2023b, 2024) require retraining the model using cross-architecture instruction embeddings. Furthermore, translating code from one ISA to another assists human analysts in understanding code from unfamiliar ISAs, supporting broader applications in code comprehension.
InnerEye Zuo et al. (2018) applies neural machine translation (NMT) techniques for binary code similarity comparison but does not perform binary code translation across ISAs. It uses two encoders from NMT models, where each encoder generates an embedding for a piece of binary code of a given pairs, and measures similarity based on embedding distance. In contrast, our approach focuses on translating binary code across ISAs.
Intermediate Representation (IR) is an architecture-agnostic representation that abstracts away various architectural differences among ISAs. However, despite this abstraction, significant variations still exist across IR code from different ISAs. Consequently, even when binaries from different ISAs—compiled from the same source code—are converted into a common IR, the resulting IR code often differs significantly, as discussed in Ahmad and Luo (2023). We compare malware detection performance across ISAs by translating binaries using FlowMalTrans against an approach that analyzes the IR code of binaries across ISAs. The results show our model enables superior cross-ISA malware detection (see the evaluation). An interesting direction is exploring translation based on IR code, which we leave for future work.